OpenStudy (anonymous):

The first term of an arithmetic series is 5 and the 22nd term is 68. Find the sum of the series

OpenStudy (anonymous):

an = a + (n-1)d use this

OpenStudy (anonymous):

find a and d?

OpenStudy (anonymous):

The first term of an arithmetic series is 5 and the 22nd term is 68. Find the sum of the series. The 22nd term (68) is the sum of the first term (5) and 21 times the common difference, d. This gives us an equation to solve for d: 5 + 21d = 68 21d = 63 d = 3 Judging by the content of these questions, you should have been given a formula for the sum of n terms of a series with first term a and common difference d. It's an + dn(n-1)/2 (although you might have a slightly different-looking formula or one with different variables). Using known values a=5, n=22, d=3 we can find the sum 5*22 + 3*22*21/2 = 110 + 33*21 = 110 + 693 = 803

OpenStudy (anonymous):

@katiebugg did u get it?

OpenStudy (anonymous):

yess lol